Solve the problem by using a system of equations.
A motel rents double rooms at $35 per day and single rooms at $23 per day. If 27 rooms were rented one day for a total of $849, how many rooms of each kind were rented?

Answer by stanbon(75887)

Solve the problem by using a system of equations.
A motel rents double rooms at $35 per day and single rooms at $23 per day.
If 27 rooms were rented one day for a total of $849, how many rooms of each kind were rented?
----
Equations:
Quantity: s + d = 27
---
Value::: 23s + 35d = 849
-----------------------------------
Multiply thru the 1st equation by 23 to get:
23s + 23d = 23*27
23s + 35d = 849
----
Subtract and solve for "d":
12d = 228
d = 19 (# of double rooms rented)
---
solve for "s":
s + d = 27
s + 19 = 27
s = 8 (# of single rooms rented)
===================
Cheers,
stan H.